Shift report 20010823 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Aug 23 2001 - 00:13:57 EDT

  • Next message: Flemming Videbaek: "Thursday meeting"

    ------------------------------------------------------------
    Date: Thu Aug 23 00:13:57 2001, Shift: 16:00-24:00
    Supervisor: JJG, with: AM,EJ
    Log book#8 pages 2 - 4
    
    Shift Summary:
    ------------------------------------------------------------
    1. Major Chaos at start of shift. No beam, No OPUS.
    2. Opus rebooted, but connection to RCF not possible due to 
    bad high throughput network card. Thus write to HPSS not 
    possible. Write to spool instead. MUST FIX TOMORROW.
    3. Runs 4341 and 4342 are pedestals runs.
    4. Beam finally came on at 22.55. T5 tripped. Run 4343. This
    is first beam since sunday night. 
    5. Raised voltage on T2A anode to 1250, since T2 has low 
    max ADC in clusters. Run 4344 in progress with 15 10**9
    ions in yellow and 4 10**9 ions in yellow. 
    All detectors appear to be operational.
    6. Have voted to end store, believing that the others 
    want to keep it. Will show our agressive attitude to data 
    taking. Is like playing poker.
    7. Tomorrow is maintenance day. Access until 4pm.
    goodnight
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Aug 23 2001 - 00:14:37 EDT